Ten yrs ago, the age of father was 33, his present age is 5 more than his wife's age; their son's age is half of his mother, the sum of their ages now os 120. Give the age of each of them
Answer by ankor@dixie-net.com(22740)

Let f = father's present age
let w = wife's age
let s = son's age
:
Write an equation for each statement
:
"Ten yrs ago, the age of father was 33,
f = 33 + 10
f = 43
:
" his present age is 5 more than his wife's age;"
f = w + 5
:
" their son's age is half of his mother,"
s = .5w
:
"the sum of their ages now is 120."
f + w + s = 120
Replace f with (w+5), replace s with .5w
(w+5) + w + .5w = 120
2.5w = 120 - 5
2.5w = 115
w = 115/2.5
w = 46 yrs is the wife's age
then
.5(46) = 23 yrs is the son's age
and
46 + 5 = 51 yrs is the father's age
;
All this was done ignoring the 1st statement which contradicts all this